Skip to Search


Support Services Meeting Notes
February 20, 2006

Present: Bill Covey, Will Chaney, Brian Keith, Barbara Oliver, Rob Roberts
  1. Howey-in-the-Hills - Junior Administrators conference. Guests include the President and the Provost. Formal business attire required. Fees are nominal. It is being held sometime this spring: April-May. Brian will forward the email to the qualifying participants.
  1. Working with minimal staff today - everyone has been farmed out on jobs so Rob had to close the office to come to this meeting. Brian is out with a death in the family. Bill is out doing property survey. Russ is helping Brad. Rob informed everyone that he is reachable via cell phone.
  2. Motor control for the air tunnel has been installed and is in use.
  3. Overhead soot in Smathers air ducts - Physical Plant did show up in Systems on Friday. Bill says if we don't start getting some action, we will smear the soot on blank, white paper and pass it around - make it public. Rob will try to contact Lacy and find out the outcome of their visit on Friday.
  1. Map server - NT13 - there was trouble getting to the server, so Joe Aufmuth rebuilt it. No further problems were reported to Will. Apparently Stephanie in the DLC was having troubles accessing it and she has somehow gotten the Provost's Office involved and offering us money to get it fixed. Now it appears to be a major fire that has to be fixed ASAP. Will met last Friday with Joe, Mark Sullivan and Winston to discuss the matter. Joe presented what the basic recommendations are to run a GIS server. Afterwards, they went down to the server and Will found the C:drive very fragmented. The other drive was not. Will has set a task for it to defrag every night. Bill suggests installing a trial version of Exec Suite software and to check the swap partition. We need to experiment more on it before we go ahead with a new purchase. The defrag has it functioning at a higher level for the moment.
  2. ALEPH print daemons - CNS changed the SMPT server name from to without informing us.
  3. The Net Agent folks are ready to stop using it. After they back up what is necessary, we will turn it off. They will do their own mail routing. Need to make sure they know their options.
  4. Public workstation renaming: PWSID. It is making it much easier to identify public workstation. The renaming has been occurring in conjunction with the WEBCT/Vista changes. Sun Java is being used.
Business Services
  1. Barbara was out all last week.
  2. Dale mentioned to Barbara the possibility of moving Judaica. Bill wants to get Dale, the Directors, HSSS & Access Services together to discuss this matter.
  3. Barbara is looking into the book mover bid.
  4. ILL problems - definition of terms is being drawn up - should assist in how thing flow.
  5. Copyright problems with ERES - Grace is following up on it.
  6. $55,000 of our money is being used to pay Preservation/Binding bills. We have set this up - changed the workflow.
  1. Frank DiTrolio has submitted his retirement. He will be here through June, but will be using vacation before then.
  2. Bob Singerman is also scheduled to retire.
  3. Barbara needs the buy-out price for both of them.
  4. Gary is leaving at the end of this month. Donation to his Diabetes Camp requested in lieu of a retirement party.
  5. Helen Jane Armstrong - emeritus status, adjunct faculty member.
  6. Computer problems in HR - Bobbie is working on it.
  7. Folks are out in HR today - some coverage issues.
  8. Three more training sessions added for Performance Evaluations. Edict from Library Council is that attendance is mandatory.
  9. Position descriptions critical to job restructuring for the Libraries are to be addressed.
  10. Staff classroom in West - have received some numbers back from Dell on the visual equipment - nothing on the laptops or carts. Dale wants a final count. Will need a copy of the numbers Dell sent to Brian.
  11. HR Web still ongoing.
  1. Friday, the mountain came to Mohammed. Meeting held with Bob Miller from Starbucks. Plans, drawings, materials were all presented for the 550 sq ft area. Most of it does not look too objectionable. Need to increase the number of seats. Meeting with the Faculty Senate this week to discuss. No cost to us or the project. May be opened 24 hours a day.
  2. Art for West wall - held up.
  3. ADA wheelchair ramp issue for the commons area. Meanwhile, $37,000 rail has gone through.
  4. Trucks have been incoming to West. Furniture starts this weekend.
  5. Carol Turner was out last week - she is back today.
  6. Move-in date still predicted as May.


University of Florida Home Page